Find 2 consecutive integral numbers whose sum 51??? I tried but i don't know what are consecutive integral numbers.
Answer by nerdybill(7384)

Find 2 consecutive integral numbers whose sum 51??? I tried but i don't know what are consecutive integral numbers.
.
"Integral numbers" are "whole numbers" that is -- integers.
.
Let x = smaller consecutive number
then
x+1 = larger consecutive number
.
from: "2 consecutive integral numbers whose sum 51" we get our equation:
x + x+1 = 51
2x+1 =51
2x = 50
x = 25 (smaller number)
.
Lareger number:
x+1 = 25+1 = 26
.
the two numbers are then 25 and 26